The cost of the packet of sugar is x cedis and the cost a tin of milk is y cedis. If 3 packets of sugar and 4 tins of milk cost 635 cedis and 4 packets of sugar and 3 tins of milk cost 695 cedis, write down an equations connecting x and y and hence find x and y
Answer by addingup(3677)

3x + 4y = 635 multiply all times 3
and
4x + 3y = 695 multiply all times -4 and add
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
9x + 12y = 1905
+
-16x + -12y = -2780
---------------------
-7x + 0y = -875
the value of y is now 0, so let's rewrite:
-7x = -875 divide both sides by 7, and remember -/- = +
x = 125 the packet of sugar is 125 cedis. Now let's plug this into one of the original equations to find y. Any of the two will work, let's use the first one:
3(125) + 4y = 635
375 + 4y = 635
4y = 260
y = 65 the tin of milk is 65 cedis
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Let's check our answers, they should solve both original equations:
3x + 4y = 635
3(125) + 4(65) = 635
375 + 260 = 635 Correct
next:
4x + 3y = 695
4(125) + 3(65) = 695
500 + 195 = 695 Also correct
:
Happy learning!